one boy can deliver newspapers on his route in 1 1/4 hours. Another boy who takes his place one day takes 15 minutes longer to deliver these papers. How long would it take to deliver the papers if the two boys worked together?
Answer by stanbon(75887)

one boy can deliver newspapers on his route in 1 1/4 hours. Another boy who takes his place one day takes 15 minutes longer to deliver these papers. How long would it take to deliver the papers if the two boys worked together?
----------
1st boy DATA:
time = 5/4 hrs/job ; rate = 4/5 job/hrs
-----
2ns boy DATA:
time = 6/4 hrs/job ; rate = 4/6 job/hrs
------
Together time = x hrs/job ; rate = 1/x job/hr
=================================
Equation:
rate + rate = together rate
4/5 + 4/6 = 1/x
(24+20)/30 = 1/x
---
x = 30/(44) = 0.6818 hrs ~ 41 minutes (time to do the job together)
=========
Cheers,
Stan H.
=========